a watertank empty in 42 seconds from 1st hole and empty in 30 seconds from the 2nd hole.in how many time the watertank will be empty when both holes are opened.make alinear equation and draw agraph.
Answer by nerdybill(7384)

a watertank empty in 42 seconds from 1st hole and empty in 30 seconds from the 2nd hole.in how many time the watertank will be empty when both holes are opened
.
Let x = time (seconds) for both holes to empty tank
then
x(1/42 + 1/30) = 1
multiply both sides by (42)(30)
x(30 + 42) = (42)(30)
x(72) = (42)(30)
x = (42)(30)/(72)
x = 17.5 seconds